4. Valle de Vallibierna

IBONES DE VALLIBIERNA (LITTLE LAKES OF VALLIBIERNA) 1 h. 30 min. (to the top)

Same route as in Ibón de Llosás. Once in the Plleta de Llosás, follow the red and white marks. After approx. 46 minutes, cross the Vallibierna ravine. Follow it on its right side and after 1 hour you will get to the lower end of the Ibón de Vallibierna (2,440 m. high). Circumvent it on its left bank. 15 minutes later you will get to the Ibón alto (upper little lake), which has another small ibón next to it.

If you feel like it, you can climb the left slope of the ibón to the hill that separates the ibones of Vallibierna and Llosás. You will need another half hour to go up to the hill (2,670 m. high) and then you can go down again by following the route of the Ibón de Llosás.